I recently found a real-life advertisement in the newspaper. (only the phone has been changed). Suppose that you have won a $20,000,000 lottery paid in 20 installments. How much would be a fair price to be paid today for the assignment of this prize? Assume the money could be invested at 6%. (Assume the lottory pays out an ordinary annuity. Round your answer to the nearest cent.
Answer #1

$20,000,000 20 4 Lotteryv amount Number of Payments Annual Payment Interest rate $1,000,000 (Assumed) 5% Fair Value Today -Present Value of future payments -1,000,000( P/A,5%, 20) $12,462,210.34 D6*PV(D7,D5,-1,0) 10 12 Hence Fair Value of the lottery today is $12,462,210.34 14

Formula sheet

A B C D E F G
2
3
4 Lottery amount 20000000
5 Number of Payments 20
6 Annual Payment 1000000 (Assumed)
7 Interest rate 0.05
8
9 Fair Value Today =Present Value of future payments
10 =1,000,000*(P/A,5%,20)
11 =D6*PV(D7,D5,-1,0) =D6*PV(D7,D5,-1,0)
12
13 Hence Fair Value of the lottery today is =D11
14